# Score Distillation via Reparametrized DDIM

### v2

@@ 1 Introduction
As an alternative to engineering-based improvements to SDS, in this paper we re-analyze the vanilla SDS algorithm to understand the underlying source of artifacts. Our key insight is that the SDS update rule steps along an approximation of the DDIM velocity field. In particular, we derive Score Distillation from DDIM with a change of variables to the space of single-step denoised images. In this light, SDS updates are nearly identical to DDIM updates, apart from one difference: while DDIM samples noise conditionally on the previous predictions, SDS resamples noise i.i.d. in every iteration. This breaks the denoising trajectory for each independent view and introduces excessive variance. Our perspective unifies DDIM and SDS and helps explain why SDS can produce blurry and over-saturated results: the variance-boosting effect of noisy guidance is usually mitigated with high Classifier-Free Guidance (CFG) [ho2022classifier] to reduce sample diversity at the cost of over-saturation [lin2024common].
@@ 2 Related work
Other papers propose heuristics that improve SDS. For example, [katzir2023noise, yu2023text, tang2023stable] decompose the guidance terms and speculate about their relative importance. Empirically, visual quality can be improved by suppressing the denoising term with negative prompts [katzir2023noise] or highlighting the classification term [yu2023text]. [lin2023magic3d, tang2023make, wang2024prolificdreamer, chen2024it3d] use multi-stage optimization: they first train a volumetric representation and then extract a mesh or voxel grid to fine-tune geometry and texture. HiFA [zhu2023hifa] uses a combination of ad-hoc techniques: additional supervision in latent space, time annealing, kernel smoothing, $z$-variance regularization, and a pretrained mono-depth estimation network to improve the quality of single-stage NeRF generation.
LucidDreamer or Interval Score Matching (ISM), hypothesizes based on empirical evidence that in SDS, the high-variance random noise term, together with the big reconstruction error of a single-step prediction cause over-smoothing [liang2023luciddreamer]. Based on these observations, the authors replace the random noise term in SDS with the noise obtained by running DDIM inversion process, and introduce multi-step denoising to improve the reconstruction quality. As we will show in section 4, the update rule of ISM is closely related to ours and can be seen as a special case of our formulation. Moreover, our analysis reveals that the added noise should be inferred conditionally on the text prompt $y$, which further leads to quality improvement.
@@ 4 Linking SDS to DDIM
In words, $x_{0}(t)$ is the noisy image at time $t$ denoised with a single step of noise prediction. Empirically, the evolution of $x_{0}(t)$ is similar to the evolution of $g(\psi,c)$—from blurry to sharp. In the left side of fig. 4 we compare these processes. This similarity motivates us to rewrite eq. 1 in terms of $x_{0}(t)$, and to understand SDS as applying similar updates to the renderings of its 3D representation.
@@
ISM as a special case. The main update formula in Interval Score Matching (ISM) [liang2023luciddreamer] is a particular case of eq. 10, where $\kappa^{t}_{y}$ is obtained via DDIM inversion without conditioning on the text prompt $y$. As demonstrated in section 5, DDIM inversion approximates a solution of eq. 10, which explains the improved performance of ISM. However, our analysis suggests that an even better-performing $\kappa^{t}_{y}$ should incorporate the text prompt $y$, which further improves the results and allows to avoid over-saturation.
@@ 6 Experiments
We demonstrate the high-fidelity 3D shapes generated with our algorithm in fig. 2 and provide more examples of $360^{\circ}$ views in section F.2.
Qualitative comparisons. Figure 7 compares 3D generation quality with the results reported in past work using a similar protocol to [wang2024prolificdreamer, katzir2023noise]. For the baselines we chose: Dreamfusion [poole2022dreamfusion] (the work we build on), Noise Free Score Distillation (NFSD) [katzir2023noise] (uses negative prompts in SDS), ProlificDreamer [wang2024prolificdreamer] (fine-tunes and trains a neural network to denoise the 3D shape), Interval Score Matching (ISM) [liang2023luciddreamer] (obtain the noise sample by inverting DDIM and perform multi-step denoising), and HiFA [zhu2023hifa] (guides in both image and latent spaces, regularizes the NeRF, and supervises the geometry with mono-depth estimation). Our figures indicate that Score Distillation via Inversion (SDI) yields similar or better results compared with state-of-the-art. Section F.3 presents more extensive comparison.
@@
Quantitative comparison. We follow [poole2022dreamfusion, yu2023text, wang2023steindreamer] to quantitatively evaluate generation quality. Table 1 provides CLIP scores [hessel2021clipscore] to measure prompt-generation alignment, computed with torchmetrics [torchmetric] and the ViT-B/32 model [radford2021learning]. We include CLIP Image Quality Assessment (IQA) [wang2023exploring] to measure quality (“Good photo” vs. “Bad photo”), sharpness (“Sharp photo” vs. “Blurry photo”), and realism (“Real photo” vs. “Abstract photo”). For each method, we test $43$ prompts with 50 views. For multi-stage baselines, we run only the first stage for fair comparison. We report the percentage of generations that run out-of-memory or generate an empty volume as “Diverged” as well as mean run time and VRAM usage. For VRAM, we average the maximum usage of GPU memory between runs. As many baselines are not open-source, we use their implementations in threestudio [threestudio2023]. SDI outperforms SDS and matches or outperforms the quality of state-of-the-art methods, offering a simple fix to SDS without additional supervision or multi-stage training.
@@
Method CLIP Score ($\uparrow$) CLIP IQA (%) $\uparrow$ Diverged (%) $\downarrow$ Time VRAM “quality” “sharpness” “real” SDS [poole2022dreamfusion], $10k$ steps $29.81\pm 2.49$ $76\pm 6.6$ $99\pm 1.2$ $98\pm 2.4$ $18.6$ 66min 6.2GB SJC [wang2023score], $10k$ steps $30.39\pm 1.98$ $76\pm 6.4$ $99\pm 0.1$ $98\pm 1.1$ $11.6$ 13min 13.1GB VSD [wang2024prolificdreamer], $25k$ steps $33.31\pm 2.39$ $77\pm 6.7$ $98\pm 1.3$ $96\pm 4.4$ $23.2$ 334min 47.9GB ESD [wang2023taming], $25k$ steps $32.79\pm 2.15$ $77\pm 7.2$ $98\pm 1.2$ $97\pm 2.7$ $14.0$ 331min 46.8GB HIFA [zhu2023hifa], $25k$ steps $32.80\pm 2.35$ $81\pm 6.5$ $98\pm 1.5$ $97\pm 1.2$ $4.7$ 235min 46.4GB SDI(ours), $10k$ steps $33.47\pm 2.49$ $82\pm 6.3$ $98\pm 1.3$ $97\pm 1.2$ $4.7$ 119min 39.2GB
@@ 6.2 Ablations
  • Random, re-sampled: Sample $\kappa^{t}_{y}(x)$ in each new update step from $\mathcal{N}(0,I)$ ;
@@
The left side of fig. 9 compares the choices for 3D generation qualitatively; the right side plots error induced in eq. 8 (re-scaled to $x_{0}$ variable due to its ambiguity around 00). As can be seen, the regressed noise has a big impact on the final generations. Both fixed point iteration and optimization via gradient descent fail to improve the approximation of $\kappa^{t}_{y}$, while fixed point iteration diverges in 3D. On the other hand, DDIM inversion yields a reasonable approximation of $\kappa^{t}_{y}$ and significantly improves 3D generation quality. We provide visual comparison of the obtained noisy images for each baseline in appendix B.
@@
CFG for inversion. [mokady2023null, miyake2023negative] report that DDIM inversion accumulates big numerical error for CFG $\gamma>0$. Surprisingly, we find that DDIM inversion for CFG $\gamma_{\text{fwd}}>0$ can be adequately estimated by running the inversion with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}$. Figure 10 compares 3D inversion strategies qualitatively and quantitatively. Naïvely taking $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ yields the biggest numerical error, while other strategies perform on par. 3D generations, illustrated on the right, show that $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ introduces excessive numerical errors, causing generation to drift in a random direction. The best parameters (as we demonstrate in appendix C) for 2D inversion ($\gamma_{\text{inv}}=\gamma_{\text{fwd}}=0$) fail to converge in 3D as there is not enough guidance toward a class sample. Introducing guidance only on the forward pass with $\gamma_{\text{inv}}=0,\gamma_{\text{fwd}}=7.5$ solves the problem, and the algorithm generates the desired 3D shape, but constantly adding CFG on each step over-saturates the image. Note, in that configuration the inversion is performed unconditionally from the text prompt, matching ISM [liang2023luciddreamer]. As we can see, prompt conditioning is an important component in eq. 8, and using $\gamma_{\text{inv}}=-7.5,\gamma_{\text{fwd}}=7.5$ cancels the over-saturation and produces accurate 3D generations.
@@ 7 Conclusion, Limitations, and Future Work
Some limitations of our algorithm motivate future work. While we improve the sample quality of each view, 3D consistency between views remains challenging; as a result, despite the convexity loss, our algorithm occasionally produces flat or concave “billboards.” A possible resolution might involve supervision with pre-trained depth or normal estimators. A related problem involves content drift from one view to another. Since there was no 3D supervision, there is little to no communication between opposite views, which can lead to inconsistent 3D assets. Stronger view conditioning, multi-view supervision, or video generation models might resolve this problem. Finally, score distillation is capped by the performance of the underlying diffusion model and is hence prone to reproduce similar “hallucinations” (e.g., text and limbs anomalies); the algorithm inherits the biases of the 2D diffusion model and can produce skewed distributions. Section F.1 demonstrates typical failure cases.
@@ Acknowledgments
The authors would like to thank Lingxiao Li and Chenyang Yuan for their thoughtful insights and feedback.
@@
The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, of Air Force Office of Scientific Research award FA9550-19-1-031, of National Science Foundation grant CHS-1955697, from the CSAIL Systems that Learn program, from the MIT–IBM Watson AI Laboratory, from the Toyota–CSAIL Joint Research Center, from a gift from Adobe Systems, and from a Google Research Scholar award.
@@
$g:\mathbb{R}^{N}\times\mathbb{C}\rightarrow\mathbb{R}^{n\times n}$ differentiable renderer
@@
$g:\mathbb{R}^{N}\times\mathbb{C}\rightarrow\mathbb{R}^{n\times n}$ differentiable renderer
@@
Section F.1 demonstrates examples of typical failure cases, including the Janus problem and the hollow face illusion.
@@
We implement our algorithm in threestudio [threestudio2023] on top of SDS [poole2022dreamfusion]. We use Stable Diffusion 2.1 [Rombach_2022_CVPR] as the diffusion model. For volumetric representation we use InstantNGP [muller2022instant]. Instead of randomly sampling time $t$ as in SDS, we maintain a global parameter $t$ that linearly decays from $1$ to $0.2$ (lower time values do not have a significant contribution). Next, for each step we render a $512\times 512$ random view and infer $\kappa^{t+\tau}$ by running DDIM inversion for $\texttt{int}(10t)$ steps - i.e. we use 10 steps of DDIM inversion for $t=1$ and linearly decrease it for smaller $t$. We use NVIDIA A6000 GPUs and run each generation for $10k$ steps with learning rate of $10^{-2}$, which takes approximately 2 wall-clock hours per shape generation.
@@ A.4 Prompts used in the quantitative evaluation
    "A DSLR photo of a beautifully decorated cupcake with sparkling sugar on top"
@@
Appendix B Visual comparison of noise patterns

### v3

@@ 1 Introduction
As an alternative to engineering-based improvements to SDS, in this paper we reanalyze the vanilla SDS algorithm to understand the underlying source of artifacts. Our key insight is that the SDS update rule steps along an approximation of the DDIM velocity field. In particular, we derive Score Distillation from DDIM with a change of variables to the space of single-step denoised images. In this light, SDS updates are nearly identical to DDIM updates, apart from one difference: while DDIM samples noise conditionally on the previous predictions, SDS resamples noise i.i.d. in every iteration. This breaks the denoising trajectory for each independent view and introduces excessive variance. Our perspective unifies DDIM and SDS and helps explain why SDS can produce blurry and over-saturated results: the variance-boosting effect of noisy guidance is usually mitigated with high Classifier-Free Guidance (CFG) [ho2022classifier] to reduce sample diversity at the cost of over-saturation [lin2024common].
@@ 2 Related work
Other papers propose heuristics that improve SDS. [katzir2023noise, yu2023text, tang2023stable] decompose the guidance terms and speculate about their relative importance. Empirically, visual quality can be improved by suppressing the denoising term with negative prompts [katzir2023noise] or highlighting the classification term [yu2023text]. [lin2023magic3d, tang2023make, wang2024prolificdreamer, chen2024it3d] use multi-stage optimization: they first train a volumetric representation and then extract a mesh or voxel grid to fine-tune geometry and texture. HiFA [zhu2023hifa] combines ad-hoc techniques: additional supervision in latent space, time annealing, kernel smoothing, $z$-variance regularization, and a pretrained mono-depth estimation network to improve the quality of single-stage NeRF generation.
LucidDreamer, or Interval Score Matching (ISM), hypothesizes based on empirical evidence that in SDS, the high-variance random noise term and large reconstruction error of a single-step prediction togehter cause over-smoothing [liang2023luciddreamer]. Based on these observations, the authors replace the random noise term in SDS with noise obtained by running DDIM inversion and introduce multi-step denoising to improve reconstruction quality. As we will show in section 4, the update rule of ISM can be seen as a special case of our formulation. Moreover, our analysis reveals that the added noise should be inferred conditionally on the text prompt $y$, which further improves quality.
@@ 4 Linking SDS to DDIM
In words, $x_{0}(t)$ is the noisy image at time $t$ denoised with a single step of noise prediction. Empirically, the evolution of $x_{0}(t)$ is similar to the evolution of $g(\psi,c)$—from blurry to sharp. The left side of fig. 4 compares these processes. This similarity motivates us to rewrite eq. 1 in terms of $x_{0}(t)$, and to understand SDS as applying similar updates to the renderings of its 3D representation.
@@
ISM as a special case. The main update formula in Interval Score Matching (ISM) [liang2023luciddreamer] is a particular case of eq. 10, where $\kappa^{t}_{y}$ is obtained via DDIM inversion without conditioning on the text prompt $y$. As demonstrated in section 5, DDIM inversion approximates a solution of eq. 10, explaining the improved performance of ISM. However, our analysis suggests that an even better-performing $\kappa^{t}_{y}$ should incorporate the text prompt $y$, which further improves the results and avoids over-saturation.
@@ 6 Experiments
We demonstrate the high-fidelity 3D shapes generated with our algorithm in fig. 2 and provide more examples of $360^{\circ}$ views in section H.2. A more detailed qualitative and quantitative comparison of our method with ISM [liang2023luciddreamer] is provided in appendix B. Additionally, we report the diversity of the generated shapes in appendix C.
Qualitative comparisons. Figure 7 compares 3D generation quality with the results reported in past work using a similar protocol to [wang2024prolificdreamer, katzir2023noise]. For the baselines we chose: Dreamfusion [poole2022dreamfusion] (the work we build on), Noise Free Score Distillation (NFSD) [katzir2023noise] (uses negative prompts in SDS), ProlificDreamer [wang2024prolificdreamer] (fine-tunes and trains a neural network to denoise the 3D shape), Interval Score Matching (ISM) [liang2023luciddreamer] (obtain the noise sample by inverting DDIM and perform multi-step denoising), and HiFA [zhu2023hifa] (guides in both image and latent spaces, regularizes the NeRF, and supervises the geometry with mono-depth estimation). Our figures indicate that Score Distillation via Inversion (SDI) yields similar or better results compared with state-of-the-art. Section H.3 presents more extensive comparison.
@@
Quantitative comparison. We follow [poole2022dreamfusion, yu2023text, wang2023steindreamer] to quantitatively evaluate generation quality. Table 1 provides CLIP scores [hessel2021clipscore] to measure prompt-generation alignment, computed with torchmetrics [torchmetric] and the ViT-B/32 model [radford2021learning]. We also report ImageReward (IR) [xu2024imagereward] to imitate possible human preference. We include CLIP Image Quality Assessment (IQA) [wang2023exploring] to measure quality (“Good photo” vs. “Bad photo”), sharpness (“Sharp photo” vs. “Blurry photo”), and realism (“Real photo” vs. “Abstract photo”). For each method, we test $43$ prompts with 50 views. For multi-stage baselines, we run only the first stage for fair comparison. We report the percentage of generations that run out-of-memory or generate an empty volume as diverged (“Div.” in the table). as well as mean run time and VRAM usage. For VRAM, we average the maximum usage of GPU memory between runs. As many baselines are not open-source, we use their implementations in threestudio [threestudio2023]. SDI outperforms SDS and matches or outperforms the quality of state-of-the-art methods, offering a simple fix to SDS without additional supervision or multi-stage training.
@@
Method CLIP Score ($\uparrow$) CLIP IQA (%) $\uparrow$ IR ($\uparrow$) Div. (%) $\downarrow$ Time VRAM “quality” “sharpness” “real” SDS [poole2022dreamfusion], $10k$ steps $29.81\pm 2.49$ $76\pm 6.6$ $99\pm 1.2$ $98\pm 2.4$ $-1.51\pm 0.83$ $18.6$ 66min 6.2GB SJC [wang2023score], $10k$ steps $30.39\pm 1.98$ $76\pm 6.4$ $99\pm 0.1$ $98\pm 1.1$ $-1.76\pm 0.51$ $11.6$ 13min 13.1GB VSD [wang2024prolificdreamer], $25k$ steps $33.31\pm 2.39$ $77\pm 6.7$ $98\pm 1.3$ $96\pm 4.4$ $-1.17\pm 0.58$ $23.2$ 334min 47.9GB ESD [wang2023taming], $25k$ steps $32.79\pm 2.15$ $77\pm 7.2$ $98\pm 1.2$ $97\pm 2.7$ $-1.20\pm 0.64$ $14.0$ 331min 46.8GB HIFA [zhu2023hifa], $25k$ steps $32.80\pm 2.35$ $81\pm 6.5$ $98\pm 1.5$ $97\pm 1.2$ $-1.16\pm 0.69$ $4.7$ 235min 46.4GB SDI(ours), $10k$ steps $33.47\pm 2.49$ $82\pm 6.3$ $98\pm 1.3$ $97\pm 1.2$ $-1.18\pm 0.59$ $4.7$ 119min 39.2GB
@@ 6.2 Ablations
  • Random, resampled: Sample $\kappa^{t}_{y}(x)$ in each new update step from $\mathcal{N}(0,I)$ ;
@@
The left side of fig. 9 compares the choices for 3D generation qualitatively; the right side plots error induced in eq. 8 (rescaled to $x_{0}$ variable due to its ambiguity around 00). As can be seen, the regressed noise has a big impact on the final generations. Both fixed point iteration and optimization via gradient descent fail to improve the approximation of $\kappa^{t}_{y}$, while fixed point iteration diverges in 3D. On the other hand, DDIM inversion yields a reasonable approximation of $\kappa^{t}_{y}$ and significantly improves 3D generation quality. We provide visual comparison of the obtained noisy images for each baseline in appendix D.
@@
CFG for inversion. [mokady2023null, miyake2023negative] report that DDIM inversion accumulates big numerical error for CFG $\gamma>0$. Surprisingly, we find that DDIM inversion for CFG $\gamma_{\text{fwd}}>0$ can be adequately estimated by running the inversion with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}$. Figure 10 compares 3D inversion strategies qualitatively and quantitatively. Naïvely taking $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ yields the biggest numerical error, while other strategies perform on par. 3D generations, illustrated on the right, show that $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ introduces excessive numerical errors, causing generation to drift in a random direction. The best parameters (as we demonstrate in appendix E) for 2D inversion ($\gamma_{\text{inv}}=\gamma_{\text{fwd}}=0$) fail to converge in 3D as there is not enough guidance toward a class sample. Introducing guidance only on the forward pass with $\gamma_{\text{inv}}=0,\gamma_{\text{fwd}}=7.5$ solves the problem, and the algorithm generates the desired 3D shape, but constantly adding CFG on each step over-saturates the image. Note in that configuration the inversion is performed unconditionally from the text prompt, matching ISM [liang2023luciddreamer]. As we can see, prompt conditioning is an important component in eq. 8, and using $\gamma_{\text{inv}}=-7.5,\gamma_{\text{fwd}}=7.5$ cancels the over-saturation and produces accurate 3D generations.
@@ 7 Conclusion, Limitations, and Future Work
Some limitations of our algorithm motivate future work. While we improve the sample quality of each view, 3D consistency between views remains challenging; as a result, despite the convexity loss, our algorithm occasionally produces flat or concave “billboards.” A possible resolution might involve supervision with pre-trained depth or normal estimators. A related problem involves content drift from one view to another. Since there was no 3D supervision, there is little to no communication between opposite views, which can lead to inconsistent 3D assets. Stronger view conditioning, multi-view supervision, or video generation models might resolve this problem. Finally, score distillation is capped by the performance of the underlying diffusion model and is hence prone to reproduce similar “hallucinations” (e.g., text and limbs anomalies); the algorithm inherits the biases of the 2D diffusion model and can produce skewed distributions. Section H.1 demonstrates typical failure cases.
@@ Acknowledgments
The authors thank Lingxiao Li and Chenyang Yuan for their thoughtful insights and feedback.
@@
The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, of National Science Foundation grant IIS-2335492, from the CSAIL Future of Data program, from the MIT–IBM Watson AI Laboratory, from the Wistron Corporation, and from the Toyota–CSAIL Joint Research Center.
@@
$g:\mathbb{R}^{N}\times\mathcal{C}\rightarrow\mathbb{R}^{n\times n}$ differentiable renderer
@@
$g:\mathbb{R}^{N}\times\mathcal{C}\rightarrow\mathbb{R}^{n\times n}$ differentiable renderer
@@
Section H.1 demonstrates examples of typical failure cases, including the Janus problem and the hollow face illusion.
@@
We implement our algorithm in threestudio [threestudio2023] on top of SDS [poole2022dreamfusion]. We use Stable Diffusion 2.1 [Rombach_2022_CVPR] as the diffusion model. For volumetric representation we use InstantNGP [muller2022instant]. Instead of randomly sampling time $t$ as in SDS, we maintain a global parameter $t$ that linearly decays from $1$ to $0.2$ (lower time values do not have a significant contribution). Next, for each step we render a $512\times 512$ random view and infer $\kappa^{t+\tau}$ by running DDIM inversion for $\texttt{int}(10t)$ steps—i.e. we use 10 steps of DDIM inversion for $t=1$ and linearly decrease it for smaller $t$. We use NVIDIA A6000 GPUs and run each generation for $10k$ steps with learning rate of $10^{-2}$, which takes approximately 2 wall clock hours per shape generation.
@@ A.4 Prompts used in the quantitative evaluation
    "A DSLR photo of a decorated cupcake with sparkling sugar on top"
@@
Appendix B Comparison with Interval Score Matching
In this section, we extensively compare our algorithm with Interval Score Matching (ISM) [liang2023luciddreamer].
Theoretical assumptions.
ISM empirically observes that “pseudo-GT” images used to guide Score Distillation Sampling (SDS) are sensitive to their input and that the single-step generation of pseudo-GT yields over-smoothing. From these observations, starting with SDS guidance, ISM adds DDIM inversion and multi-step generation to empirically improve the stability of the guidance. In contrast, our work starts with 2D diffusion sampling to re-derive score-distillation guidance and motivate improvements. That is, our work formally connects SDS to well-justified techniques in 2D sampling.
DDIM inversion.
In ISM, the empirically motivated DDIM inversion is at the basis of the derivation of the final update rule. We suggest a general form of the noise term eq. 8, for which DDIM inversion is just one possible solution. Our theoretical insights are agnostic to particular algorithms of root-finding, which makes it possible to use more efficient solutions in future research (e.g. train diffusion models as invertible maps to sample noise faster).
Guidance term.
The update rules provided in our eq. 10 and ISM’s eq.17 have two main differences:
  1. 1.
    Full vs. interval noise. Assuming DDIM inversion finds a perfect root of our stationary equation, ISM’s update rule can take a similar form to ours eq. 10 (interval noise is equal to $\kappa$ if eq. 8 is satisfied). However, as shown in fig. 9, DDIM inversion does not find a perfect root, and thus the two forms are not equivalent. Our theory shows that the full noise term is more accurate. We also show the effect of choosing one term vs. another in fig. 14.
  2. 2.
    Conditional vs. unconditional inversion. Our derivation hints that the roots of eq. 8 are prompt-dependent, motivating our use of conditional DDIM inversion (not used in ISM). Our fig. 10 shows how unconditional inversion yields over-saturation.
Figure 13: Comparison of SDI (ours, top) ISM (middle), and SDI in ISM’s code base (ours, bottom). To eliminate factors such as 3D representation and initialization, we re-implement our algorithm in the official codebase of ISM. See fig. 14 for the effect of each change made.
Practical Results.
To control for different design choices (Gaussian Splattings in ISM vs. NeRF in ours, etc.) we re-implemented our algorithm in the code base of ISM with only the minimal changes discussed in the previous section.
Figure 14: Ablation of changes made to the code base of Interval Score Matching (ISM). Prompt ”A DSLR photo of a hamburger“. We start with the official implementation of ISM and add only four changes to match ours (all the hyper-parameters remain tuned for ISM): 1. instead of sampling t from U(0.02, 0.5) we sample it from U(0.1, 0.98); 2. introduce linear time annealing; 3. use conditional DDIM inversion with negative guidance; 4. use SDI guidance, i.e. second term in the guidance difference is the full noise added to
Figure 13 provides a qualitative comparison using the prompts and settings in ISM’s code. Figure 14 shows the effect of each change made to ISM guidance. Table 2 the quantitative comparison (ISM code base for both).
Table 2: Quantitative comparisons between SDI (ours) and ISM.
Method CLIP Score ($\uparrow$) CLIP IQA (%, $\uparrow$) ImageReward ($\uparrow$) Time VRAM “quality” “sharpness” “real” ISM, 5k steps 28.60±2.03 0.85±0.02 0.98±0.01 0.98±0.01 -0.52±0.48 45min 15.4GB SDI (ours), 5k steps 28.47±1.29 0.88±0.03 0.99±0.00 0.98±0.01 -0.30±0.32 43min 15.4GB
Summary.
We bridge the gap between experimentally-based score distillation techniques and theoretically-justified 2D sampling. Both SDS and ISM can be seen as different approaches to finding roots of eq. 8. This theoretical insight allows us to modify both ISM and SDS, reducing over-saturation for the first and improving general quality of the second.
Appendix C Diversity of the generations
Next, we explore the diversity of the results generated by our method. Figure 15 depicts generated shapes for different seeds and prompts. Generations with our method exhibit a certain degree of diversity in local details, but are mostly the same at a coarse level. This is similar to the reported results in other works on score distillation [poole2022dreamfusion, wang2024prolificdreamer]. We plan to address this problem in future research.
Figure 15: Examples of generations with SDI (ours) for different seed values. Prompts are “Pumpkin head zombie, skinny, highly detailed, photorealistic” (top row), “A highly-detailed sandcastle” (middle row), “An ice cream sundae” (bottom row).
Additionally, we noticed a minimal diversity in human generations. For example, fig. 16 demonstrates generated shapes for 3 prompts related to different professions.
Figure 16: Examples of limited diversity across different human-related prompts: “a photo of a firefighter”, “a photo of a policeman”, “a photo of a soldier.”
Appendix D Visual comparison of noise patterns